OpenStudy (anonymous):

HELP me solve this! .... Mutt and Jeff need to paint a fence. Mutt can do the job alone 5 hours faster than Jeff. If together they work for 25 hours and finish only 7/8 of the job, how long would Jeff need to do the job alone?

OpenStudy (mertsj):

Jeff can do the job in x hours so he does 1/x jobs per hour Mutt can do the job in x-5 hours so he does 1/(x-5) jobs per hour. They each work for 25 hours. Jeff does this part of the job in 25 hours: 25/x Mutt does this part of the job in 25 hours: 25/(x-5) Together they do: 25/x+25/(x-5) and they get 7/8 of the job done. So:

OpenStudy (mertsj):

\[\frac{25}{x}+\frac{25}{(x-5)}=\frac{7}{8}\]

OpenStudy (mertsj):

Solve that

OpenStudy (anonymous):

i got to that part already and im getting an incorrect answer. can you please tell me what you come up with?

OpenStudy (mertsj):

60 hours for Jeff and 55 for Mutt

OpenStudy (anonymous):

60 is incorrect as well

OpenStudy (anonymous):

im bout to give up